# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ 5. Random attractors for stochastic third-grade fluids equations: Unbounded domains
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Suppose that $\boldsymbol{y}$ solves system (3.19) on the time interval $[a,\infty)$ with $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}(\mathbb{R};\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}(\mathbb{R};\mathbb{W}^{1,4}(\mathscr{O}))$, $\chi\geq 0$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. Then, for any $t\geq s\geq a,$
@@
$\left|\alpha\left<\mathscr{J}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{y}\right>\right|$ $\leq\frac{\nu(1-\varepsilon_{0})}{4}\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\frac{\beta(1-\varepsilon_{0})}{2}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}$ (5.4)
$\left|\beta\left<\mathscr{K}(\boldsymbol{y}+\boldsymbol{z}),\boldsymbol{z}\right>\right|$ $\leq\frac{\beta\varepsilon_{0}}{8}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+C\|\mathrm{A}(\boldsymbol{z})\|^{4}_{4},$ (5.5)
@@
Using (5.11), Lemmas 3.5 and 3.6, and the fact that $\lim\limits_{s\to-\infty}(1-s)e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)s}=0,$ we have that there exists a $t^{\ast}_{\mathrm{D}}(\omega)\geq t_{\mathrm{D}}(\omega)$ such that for all $s\leq-t^{\ast}_{\mathrm{D}}(\omega)$
@@
Using (5.11), Lemmas 3.5 and 3.6, and the fact that $\lim\limits_{s\to-\infty}\left\{-se^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\right\}=0,$ we have that there exists a $t^{\ast\ast}_{\mathrm{D}}(\omega)\geq t_{\mathrm{D}}(\omega)$ such that for all $s\leq-t^{\ast\ast}_{\mathrm{D}}(\omega)$
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied and let us denote by $\boldsymbol{y}(t,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}),$ the solution of system (3.19). Assume that $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\{\boldsymbol{x}_{m}\}_{m\in\mathbb{N}}$ be a bounded sequence in $\mathbb{H}$. Then, there exists $\tilde{\boldsymbol{y}}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ such that along a subsequence
$\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}_{m})\ \to\tilde{\boldsymbol{y}}\ \text{ in }\ \mathrm{L}^{2}(0,T;\mathbb{L}^{2}_{\mathrm{loc}}(\mathscr{O})),$ (5.16)
@@
such that, as $m^{\prime}\to\infty,$ (by the Banach-Alaoglu theorem)
[…]
@@
Integrating from 00 to $T-a$, we further find
@@
$\lim_{a\to 0}\sup_{m}\int_{0}^{T-a}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{2}^{2}\mathrm{d}t=0.$ (5.22)
@@
$\lim_{a\to 0}\sup_{m}\int_{0}^{T-a}\|\boldsymbol{y}_{m,k}(t+a)-\boldsymbol{y}_{m,k}(t)\|_{\mathbb{L}^{2}(\mathscr{O}_{2k})}^{2}\mathrm{d}t=0,$ (5.23)
@@
Since the injection $\mathbb{H}_{0}^{1}(\mathscr{O}_{2k})\subset\mathbb{L}^{2}{(\mathscr{O}_{2k})}$ is compact, we can apply [50, Theorem 16.3] (see [63, Theorem 13.3] also) to obtain
@@
$\boldsymbol{y}_{m}\to\tilde{\boldsymbol{y}}\ \text{ in }\mathrm{L}^{2}(0,T;\mathbb{L}^{2}(\mathscr{O}_{k})),$
@@
Taking the divergence of the first equation of (3.18), we obtain
$-\Delta\widehat{\textbf{P}}$
$=\nabla\cdot\left[\big{(}(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla\big{)}(\boldsymbol{y}+\boldsymbol{z})+\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})+\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))\right]-\nabla\cdot\boldsymbol{f}$
=[$=\nabla\cdot\left[\nabla\cdot\left\{\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}+\alpha(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]+\beta\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\right\}\right]-\nabla\cdot\boldsymbol{f}$
$=:-\Delta\widehat{\textbf{P}}_{1}-\Delta\widehat{\textbf{P}}_{2},$
in the weak sense, where
$-\Delta\widehat{\textbf{P}}_{1}$ =[$=\nabla\cdot\left[\nabla\cdot\left\{\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}+\alpha(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]-\nabla\cdot\boldsymbol{f},$
$-\Delta\widehat{\textbf{P}}_{2}$ $=\beta\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\right\}\right].$
Note that, by Lemma 5.2, we have for $s\leq-\mathscr{T}=\max\{t_{\mathrm{D}}^{\ast}(\omega),t_{\mathrm{D}}^{\ast\ast}(\omega)\}$
$e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\bigg{(}\nabla\cdot\left[\nabla\cdot\left\{\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}+\alpha(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]-\nabla\cdot\boldsymbol{f}\bigg{)}$ $\in\mathrm{L}^{2}(s,0;\mathbb{H}^{-2}(\mathscr{O}))$
and
$e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\bigg{(}\beta\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v})\right\}\right]\bigg{)}$ $\in\mathrm{L}^{\frac{4}{3}}(s,0;\mathbb{W}^{-2,\frac{4}{3}}(\mathscr{O}))$
which gives
$e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\widehat{\textbf{P}}_{1}\in\mathrm{L}^{2}(s,0;\mathbb{L}^{2}(\mathscr{O}))\;\text{ and }\;e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\widehat{\textbf{P}}_{2}\in\mathrm{L}^{\frac{4}{3}}(s,0;\mathbb{L}^{\frac{4}{3}}(\mathscr{O})).$
Also note that an application of Poincaré inequality (1.8) allow us to write
@@
Taking the inner product of the first equation of (3.18) with $\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}$ in $\mathbb{L}^{2}(\mathscr{O})$, we have
@@
$=\underbrace{\nu\int_{\mathscr{O}}(\Delta\boldsymbol{y})\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{1}(k,t)}-\underbrace{b\left(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y}+\boldsymbol{z},\Lambda^{2}\left(|x|^{2}k^{-2}\right)(\boldsymbol{y}+\boldsymbol{z})\right)}_{I_{2}(k,t)}+\underbrace{b\left(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y}+\boldsymbol{z},\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{z}\right)}_{I_{3}(k,t)}$
$\quad+\underbrace{\alpha\int_{\mathscr{O}}(\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}))\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{4}(k,t)}+\underbrace{\beta\int_{\mathscr{O}}(\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})))\Lambda^{2}\left(|x|^{2}k^{-2}\right)(\boldsymbol{y}+\boldsymbol{z})\mathrm{d}x}_{I_{5}(k,t)}$
$\quad-\underbrace{\beta\int_{\mathscr{O}}(\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})))\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{z}\mathrm{d}x}_{I_{6}(k,t)}-\underbrace{\int_{\mathscr{O}}(\nabla\widehat{\textbf{P}})\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\mathrm{d}x}_{I_{7}(k,t)}$
@@
Let us now estimate each term on the right hand side of (5). Integration by parts, (1.4) (see Remark 1.1 also), and Hölder’s, Gagliardo-Nirenberg’s (2.1) and Young’s inequalities help us to obtain
@@
$\leq-\frac{\nu}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x+\frac{4\sqrt{2}\nu}{k}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}\left|\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)\right|\left|\boldsymbol{y}\right|\left|\nabla\boldsymbol{y}\right|\mathrm{d}x$
@@
$-I_{2}(k,t)$ $=\left|\frac{2}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)x\cdot(\boldsymbol{y}+\boldsymbol{z})|\boldsymbol{y}+\boldsymbol{z}|^{2}\mathrm{d}x\right|$
@@
$I_{3}(k,t)$ $\leq\int_{\mathscr{O}}|\boldsymbol{y}+\boldsymbol{z}||\nabla(\boldsymbol{y}+\boldsymbol{z})|\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|\mathrm{d}x$
@@
$I_{4}(k,t)$ $\leq\left|\frac{\alpha}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)[(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}:\mathrm{A}(\boldsymbol{y})]\mathrm{d}x\right|$
@@
$-I_{6}(k,t)$ $\leq\beta\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|\nabla\boldsymbol{z}|\mathrm{d}x+\frac{4\beta}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|x||\boldsymbol{z}|\mathrm{d}x$
@@
$-I_{7}(k,t)$ $\leq\left|\frac{4}{k^{2}}\int_{\mathscr{O}}\widehat{\textbf{P}}\;\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)(x\cdot\boldsymbol{y})\mathrm{d}x\right|$
$\leq\frac{C}{k}\int\limits_{\mathscr{O}\cap\{k\leq|x|\leq\sqrt{2}k\}}\left|\widehat{\textbf{P}}\right|\left|\boldsymbol{y}\right|\mathrm{d}x$
$\leq\frac{C}{k}\int_{\mathscr{O}}\bigg{\{}\left|(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\big{(}(\boldsymbol{y}+\boldsymbol{z})\otimes(\boldsymbol{y}+\boldsymbol{z})\big{)}\right]\right|+\left|(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}\right\}\right]\right|$
$\qquad\qquad+\left|(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\right\}\right]\right|+\left|(-\Delta)^{-1}\left[\nabla\cdot\boldsymbol{f}\right]\right|\bigg{\}}\left|\boldsymbol{y}\right|\mathrm{d}x$
$\leq\frac{C}{k}\left[\|\boldsymbol{y}+\boldsymbol{z}\|^{2}_{4}\|\boldsymbol{y}\|_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{2}_{4}\|\boldsymbol{y}\|_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{3}_{4}\|\boldsymbol{y}\|_{4}+\|\boldsymbol{f}\|_{2}\|\boldsymbol{y}\|_{2}\right]$
$\leq\frac{C}{k}\left[\|\boldsymbol{y}\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{4}+\|\boldsymbol{y}\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{f}\|^{2}_{2}\right]$
$\leq\frac{C}{k}\left[\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{f}\|^{2}_{2}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}\right],$ (5.37)
$I_{8}(k,t)$ $\leq\frac{\nu\lambda\varepsilon_{0}}{8}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{f}|^{2}\mathrm{d}x,$ (5.38)
$I_{9}(k,t)$ $\leq\frac{\nu\lambda\varepsilon_{0}}{8}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|^{2}\mathrm{d}x.$ (5.39)
@@
$\quad+\frac{C}{k}\left[\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}\|^{2}_{2}+\|\boldsymbol{f}\|^{2}_{2}\right],$

### v2

@@ 5. Random attractors for a subclass of stochastic third-grade fluids equations: Unbounded domains
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Suppose that $\boldsymbol{y}$ solves system (3.19) on the time interval $a,\infty)$ with $\boldsymbol{z}\in\mathrm{L}^{2}_{\mathrm{loc}}(\mathbb{R};\mathbb{H})\cap\mathrm{L}^{4}_{\mathrm{loc}}(\mathbb{R};\mathbb{W}^{1,4}(\mathscr{O}))$, $\chi\geq 0$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. Then, for any $t\geq s\geq a,$
@@
[…]
@@
Using (5.11), Lemmas 3.5 and 3.6, and the fact that $\lim\limits_{s\rightarrow-\infty}(1-s)e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)s}=0,$ we have that there exists a $t^{\ast}_{\mathrm{D}}(\omega)\geq t_{\mathrm{D}}(\omega)$ such that for all $s\leq-t^{\ast}_{\mathrm{D}}(\omega)$
@@
Using (5.11), Lemmas 3.5 and 3.6, and the fact that $\lim\limits_{s\rightarrow-\infty}\left\{-se^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)s}\right\}=0,$ we have that there exists a $t^{\ast\ast}_{\mathrm{D}}(\omega)\geq t_{\mathrm{D}}(\omega)$ such that for all $s\leq-t^{\ast\ast}_{\mathrm{D}}(\omega)$
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied and let us denote by $\boldsymbol{y}(t,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}),$ the solution of system (3.19). Assume that $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\{\boldsymbol{x}_{m}\}_{m\in\mathbb{N}}$ be a bounded sequence in $\mathbb{H}$. Then, there exists $\tilde{\boldsymbol{y}}\in\mathrm{L}^{\infty}(0,T;\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ such that along a subsequence
$\boldsymbol{y}(\cdot,\boldsymbol{z},\boldsymbol{f},\boldsymbol{x}_{m})\ \rightarrow\tilde{\boldsymbol{y}}\ \text{ in }\ \mathrm{L}^{2}(0,T;\mathbb{L}^{2}_{\mathrm{loc}}(\mathscr{O})),$ (5.16)
@@
such that, as $m^{\prime}\rightarrow\infty,$ (by the Banach-Alaoglu theorem)
$\begin{cases}\boldsymbol{y}_{m^{\prime}}\xrightharpoonup{w^{*}}\widetilde{\boldsymbol{y}}\ \text{ in }\ \mathrm{L}^{\infty}(0,T;\mathbb{H})\\ \boldsymbol{y}_{m^{\prime}}\xrightharpoonup{w}\widetilde{\boldsymbol{y}}\ \text{ in }\ \mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})).\end{cases}$ (5.18)
@@
Integrating from 0 to TaT-a, we further find
@@
$\lim_{a\rightarrow 0}\sup_{m}\int_{0}^{T-a}\|\boldsymbol{y}_{m}(t+a)-\boldsymbol{y}_{m}(t)\|_{2}^{2}\mathrm{d}t=0.$ (5.22)
@@
$\lim_{a\rightarrow 0}\sup_{m}\int_{0}^{T-a}\|\boldsymbol{y}_{m,k}(t+a)-\boldsymbol{y}_{m,k}(t)\|_{\mathbb{L}^{2}(\mathscr{O}_{2k})}^{2}\mathrm{d}t=0,$ (5.23)
@@
Since the injection $\mathbb{H}_{0}^{1}(\mathscr{O}_{2k})\subset\mathbb{L}^{2}{(\mathscr{O}_{2k})}$ is compact, we can apply [56, Theorem 16.3] (see [72, Theorem 13.3] also) to obtain
@@
$\boldsymbol{y}_{m}\rightarrow\tilde{\boldsymbol{y}}\ \text{ in }\mathrm{L}^{2}(0,T;\mathbb{L}^{2}(\mathscr{O}_{k})),$
@@
Note that an application of Poincaré inequality (1.8) allow us to write
@@
Taking the inner product of the first equation of (3.18) with $\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}$, we have
@@
$=\underbrace{\nu\left\langle\Delta\boldsymbol{y},\Lambda^{2}\left(|x|^{2}k^{-2}\right)\right\rangle}_{I_{1}(k,t)}-\underbrace{b\left(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y}+\boldsymbol{z},\Lambda^{2}\left(|x|^{2}k^{-2}\right)(\boldsymbol{y}+\boldsymbol{z})\right)}_{I_{2}(k,t)}+\underbrace{b\left(\boldsymbol{y}+\boldsymbol{z},\boldsymbol{y}+\boldsymbol{z},\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{z}\right)}_{I_{3}(k,t)}$
$\quad+\underbrace{\alpha\left\langle\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}),\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\right\rangle}_{I_{4}(k,t)}+\underbrace{\beta\left\langle\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})),\Lambda^{2}\left(|x|^{2}k^{-2}\right)(\boldsymbol{y}+\boldsymbol{z})\right\rangle}_{I_{5}(k,t)}$
$\quad-\underbrace{\beta\left\langle\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})),\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{z}\right\rangle}_{I_{6}(k,t)}-\underbrace{\left\langle\nabla\widehat{\textbf{P}},\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\right\rangle}_{I_{7}(k,t)}$
@@
Let us now estimate each term on the right hand side of (5). Integration by parts, (1.4) (see Remark 1.1 also), and Hölder’s, Gagliardo-Nirenberg’s (2.1), Korn’s (2.2) and Young’s inequalities help us to obtain
@@
[…]
@@
$|I_{2}(k,t)|$ $=\left|\frac{2}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)x\cdot(\boldsymbol{y}+\boldsymbol{z})|\boldsymbol{y}+\boldsymbol{z}|^{2}\mathrm{d}x\right|$
@@
$|I_{3}(k,t)|$ $\leq\int_{\mathscr{O}}|\boldsymbol{y}+\boldsymbol{z}||\nabla(\boldsymbol{y}+\boldsymbol{z})|\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|\mathrm{d}x$
@@
$|I_{4}(k,t)|$ $\leq\left|\frac{\alpha}{2}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)[(\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2}:\mathrm{A}(\boldsymbol{y})]\mathrm{d}x\right|$
@@
$|I_{6}(k,t)|$ $\leq\beta\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|\nabla\boldsymbol{z}|\mathrm{d}x+\frac{4\beta}{k^{2}}\int_{\mathscr{O}}\Lambda\left(|x|^{2}k^{-2}\right)\Lambda^{\prime}\left(|x|^{2}k^{-2}\right)|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{3}|x||\boldsymbol{z}|\mathrm{d}x$
@@
$|I_{7}(k,t)|$ $\leq\left|\left\langle\nabla\widehat{\textbf{P}}_{1},\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\right\rangle\right|+\left|\left\langle\nabla\widehat{\textbf{P}}_{2},\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\right\rangle\right|$
$\leq\|\nabla\widehat{\textbf{P}}_{1}\|_{\mathbb{H}^{-1}}\|\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\|_{\mathbb{H}^{1}}+\|\nabla\widehat{\textbf{P}}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}\|\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\|_{\mathbb{W}^{1,4}}$
$\leq\|\nabla\widehat{\textbf{P}}_{1}\|_{\mathbb{H}^{-1}}\bigg(\|\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\|^{2}_{2}+\|\nabla[\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}]\|^{2}_{2}\bigg)^{\frac{1}{2}}$
$\quad+\|\nabla\widehat{\textbf{P}}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg(\|\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}\|^{4}_{4}+\|\nabla[\Lambda^{2}\left(|x|^{2}k^{-2}\right)\boldsymbol{y}]\|^{4}_{4}\bigg)^{\frac{1}{4}}$
$\leq C\|\nabla\widehat{\textbf{P}}_{1}\|_{\mathbb{H}^{-1}}\bigg(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}|^{2}\mathrm{d}x+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x\bigg)^{\frac{1}{2}}$
$+C\|\nabla\widehat{\textbf{P}}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg(\bigg\{\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}|^{2}\mathrm{d}x\bigg\}^{2}+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\bigg)^{\frac{1}{4}},$ (5.37)
$|I_{8}(k,t)|$ $\leq\frac{\nu\lambda\varepsilon_{0}}{8}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{f}|^{2}\mathrm{d}x,$ (5.38)
$|I_{9}(k,t)|$ $\leq\frac{\nu\lambda\varepsilon_{0}}{8}\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{v}|^{2}\mathrm{d}x+C\int_{\mathscr{O}}\Lambda^{2}\left(|x|^{2}k^{-2}\right)|\boldsymbol{z}|^{2}\mathrm{d}x.$ (5.39)
@@
$\quad+C\|\nabla\widehat{\textbf{P}}_{1}\|_{\mathbb{H}^{-1}}\bigg(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}|^{2}\mathrm{d}x+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x\bigg)^{\frac{1}{2}}$
$\quad+C\|\nabla\widehat{\textbf{P}}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg(\bigg\{\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}|^{2}\mathrm{d}x\bigg\}^{2}+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\bigg)^{\frac{1}{4}}$
$\quad+\frac{C}{k}\left[\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}\|^{2}_{2}\right],$